The customer account at a certain department store has an average balance of 480 kwacha and the standard deviation of 160 kwacha. Assuming normal distribution, What proportion of the account is between 240 and 360.


1
Expert's answer
2021-01-21T00:17:35-0500

"P(240<X<360)=P(\frac{240-480}{160}<Z<\frac{360-480}{160})=P(-1.5<Z<-0.75)="

"=P(Z<-0.75)-P(Z<-1.5)=P(Z<1.5)-P(Z<0.75)="

"=0.9332-0.7734=0.1598."